# Picky elements, subnormalisers, and character correspondences

### v1

@@ 6.1. Subnormalisers of unipotent elements
By [24, Prop. 2.9] we may embed ${\mathbf{U}}$ into the (unipotent) group of upper uni-triangular matrices of ${\operatorname{GL}}_{n}$ for a suitable $n$, so without loss we may assume ${\mathbf{U}}$ is connected. By [24, Cor. 2.10] the unipotent group ${\mathbf{U}}$ is nilpotent, so by [19, Prop. 17.4], for any proper closed subgroup ${\mathbf{H}}<{\mathbf{U}}$ we have $\dim{\mathbf{H}}<\dim{\mathbf{N}}_{\mathbf{U}}({\mathbf{H}})$. Thus by induction on the dimension, any closed subgroup is subnormal in ${\mathbf{U}}$. As $\langle x\rangle$ is normal in its abelian (see [19, Lemma 15.1.C]) closure $\overline{\langle x\rangle}$ this achieves the proof. ∎
@@
Now the Zariski closure $\overline{{\widehat{R_{u}}}(H)}$ is unipotent [24, Prop. 2.9], giving $\langle x\rangle{\,\triangleleft\triangleleft\,}\overline{{\widehat{R_{u}}}(H)}$ by Lemma 6.1 and then of course $\langle x\rangle{\,\triangleleft\triangleleft\,}{\widehat{R_{u}}}(H)$. Hence, $\langle x\rangle{\,\triangleleft\triangleleft\,}{\mathbf{N}}_{\mathbf{G}}({\widehat{R_{u}}}(H))\geq H$ and the last assertion follows. ∎
@@
Let ${\mathbf{U}}\leq{\mathbf{G}}$ be a closed unipotent subgroup with ${\mathbf{U}}={\widehat{R_{u}}}({\mathbf{N}}_{\mathbf{G}}({\mathbf{U}}))$. By [24, Cor. 17.15], ${\mathbf{U}}$ lies in some Borel subgroup of ${\mathbf{G}}$ and then by [24, Thm 17.10] there is a parabolic subgroup ${\mathbf{P}}\leq{\mathbf{G}}$ with ${\mathbf{U}}\leq R_{u}({\mathbf{P}})$ and ${\mathbf{N}}_{\mathbf{G}}({\mathbf{U}})\leq{\mathbf{P}}$. In particular, ${\mathbf{N}}_{\mathbf{G}}({\mathbf{U}})$ normalises $R_{u}({\mathbf{P}})$. Thus, $V:={\mathbf{N}}_{R_{u}({\mathbf{P}})}({\mathbf{U}})$ is contained in and normalised by ${\mathbf{N}}_{\mathbf{G}}({\mathbf{U}})$, so contained in ${\widehat{R_{u}}}({\mathbf{N}}_{\mathbf{G}}({\mathbf{U}}))={\mathbf{U}}$, whence $V={\mathbf{U}}$. By [19, Prop. 17.4] this forces ${\mathbf{U}}=R_{u}({\mathbf{P}})$.
Conversely, if ${\mathbf{P}}\leq{\mathbf{G}}$ is parabolic, hence connected, then ${\mathbf{P}}/R_{u}({\mathbf{P}})$ is connected reductive, and so has no non-trivial normal unipotent subgroups (since any such would have to be finite, hence central, but all central elements of connected reductive groups are semisimple, e.g. by [24, Cor. 8.13(b)]), whence ${\widehat{R_{u}}}({\mathbf{P}})=R_{u}({\mathbf{P}})$. ∎
@@
By Lemma 6.1, ${\operatorname{Sub}}_{\mathbf{G}}(x)$ contains all Borel subgroups of ${\mathbf{G}}$ containing $x$. For the converse, assume $\langle x\rangle{\,\triangleleft\triangleleft\,}H$ for some subgroup $H\leq{\mathbf{G}}$, where without loss of generality $H$ is maximal with respect to inclusion. By Lemma 6.2 then $H={\mathbf{N}}_{\mathbf{G}}({\widehat{R_{u}}}(H))$ and $x\in Q:={\widehat{R_{u}}}(H)$. Now $Q$ lies in the unipotent radical ${\mathbf{U}}$ of some Borel subgroup of ${\mathbf{G}}$ ([24, Cor. 17.15]). If $Q<{\mathbf{U}}$ then $Q_{1}:={\mathbf{N}}_{\mathbf{U}}(Q)>Q$, and $\langle x\rangle{\,\triangleleft\triangleleft\,}\langle H,Q_{1}\rangle>H$, a contradiction, so in fact $Q={\mathbf{U}}$ and $H={\mathbf{N}}_{\mathbf{G}}(Q)={\mathbf{N}}_{\mathbf{G}}({\mathbf{U}})$ is closed (see [24, Ex. 10.18]). Then ${\mathbf{N}}_{H}({\mathbf{U}})^{\circ}$ is a Borel subgroup of $H$, and thus $H^{\circ}$ is generated by the $H$-conjugates of ${\mathbf{N}}_{H}({\mathbf{U}})^{\circ}$ (see [24, Thm 6.10]), all of which contain $Q$ and hence $x$. Thus, $H$ is generated by the normalisers (in $H$) of the maximal unipotent subgroups (of $H$ and hence of $G$) it contains. The claim follows. ∎
@@ 6.2. An extension to disconnected groups
In this subsection we consider a slightly different situation. Let now ${\mathbf{G}}$ be the extension of a connected reductive algebraic group ${\mathbf{G}}^{\circ}$ in characteristic $p$ by a graph automorphism $\sigma$ of order $p$. We are interested in unipotent elements in non-trivial cosets of ${\mathbf{G}}^{\circ}$ in ${\mathbf{G}}$. Every coset of ${\mathbf{G}}^{\circ}$ in ${\mathbf{G}}$ contains a unique class of regular unipotent elements satisfying analogous properties to the case of connected groups (see [33, Prop. II.10.2]).
Let ${\mathbf{B}}\leq{\mathbf{G}}$ be the normaliser in ${\mathbf{G}}$ of a $\sigma$-stable Borel subgroup ${\mathbf{B}}^{\circ}$ of ${\mathbf{G}}^{\circ}$, with $\sigma$-stable maximal torus ${\mathbf{T}}^{\circ}$ and unipotent radical ${\mathbf{U}}={\mathbf{U}}^{\circ}$. Let $W={\mathbf{N}}_{{\mathbf{G}}^{\circ}}({\mathbf{T}}^{\circ})/{\mathbf{T}}^{\circ}$ be the Weyl group of ${\mathbf{G}}^{\circ}$; it is normalised by $\sigma$, and $\sigma$ acts on its set of roots $\Phi$ with respect to ${\mathbf{T}}^{\circ}$, permuting its set $\Delta$ of simple roots. We let $\bar{\Delta}$ denote the set of $\sigma$-orbits in $\Delta$, and we write $\bar{S}$ for the set of simple reflections of $W^{\sigma}$ constructed as in [24, Lemma 23.3].
@@
Since ${\mathbf{U}}$ is a maximal unipotent subgroup in ${\mathbf{G}}^{\circ}$ and normalised by $\sigma$ we may assume $x\in{\mathbf{U}}\sigma$. So $x=x_{0}\sigma$ where $x_{0}$ is a product of root elements for ${\mathbf{G}}^{\circ}$. If root elements for representatives from all $\sigma$-orbits of simple roots occur in $x_{0}$, then $x$ is regular and hence contained in a unique Borel subgroup of ${\mathbf{G}}$ by [33, Prop. II.10.2], so ${\operatorname{Sub}}_{\mathbf{G}}(x)={\mathbf{C}}_{\mathbf{B}}(\sigma)$. Representatives for the non-regular outer unipotent classes in types $A_{2},A_{3}$ and $A_{4}$ are given in [22, Tab. 4]. It transpires that in each case at least one of the given representatives in the finite group for a class in the algebraic group has the property that its image under any reflection in $\bar{S}$ is still in ${\mathbf{U}}\sigma$, and thus $x$ lies in a unipotent normal subgroup of any ${\mathbf{P}}^{\sigma}$, where ${\mathbf{P}}$ runs over the normalisers in ${\mathbf{G}}$ of the minimal $\sigma$-stable standard parabolic subgroups of ${\mathbf{G}}^{\circ}$. Since these generate ${\mathbf{G}}$, we have ${\operatorname{Sub}}_{\mathbf{G}}(x)={\mathbf{G}}$. ∎
@@
First, by [33, Prop. II.10.2] the regular unipotent elements in ${\mathbf{G}}^{\circ}\sigma$ are picky, arguing as in the proof of Proposition 3.1. Now assume $x\in{\mathbf{G}}^{\circ}\sigma$ is unipotent non-regular. Again we may assume $x\in{\mathbf{U}}\sigma$. If ${\mathbf{C}}_{{\mathbf{G}}^{\circ}}(\sigma)$ has rank 1 or 2, we are done by Proposition 6.6, or ${\mathbf{G}}^{\circ}$ is of type $D_{4}$, $p=3$ and $\sigma$ induces triality. In the latter case, inspection of the representatives given in [22, Tab. 8] shows that $x$ can be chosen to lie in the maximal normal unipotent subgroups of ${\mathbf{C}}_{\mathbf{P}}(\sigma)$, where ${\mathbf{P}}$ runs over the normalisers in ${\mathbf{G}}$ of the minimal $\sigma$-stable standard parabolic subgroups of ${\mathbf{G}}^{\circ}$ and thus ${\operatorname{Sub}}_{\mathbf{G}}(x)={\mathbf{G}}$.
So we may now assume that $p=2$, ${\mathbf{C}}_{\mathbf{G}}^{\circ}(\sigma)$ has rank at least 3, and that the claim has been shown for all groups of smaller rank. As $x$ is not regular, by [33, Prop. II.10.2] it is a product of root elements not involving representatives from at least one orbit of simple roots $\alpha\in\bar{\Delta}$. Consider the Dynkin diagram of ${\mathbf{C}}_{{\mathbf{G}}}^{\circ}(\sigma)$, with nodes labelled by $\bar{\Delta}$. If $\alpha$ is not an end node, then we can conclude verbatim by the arguments in the proof of Proposition 3.6. If $\alpha$ is an end node, then the image of $x$ in the standard Levi subgroup corresponding to $\alpha$ and the adjacent node $\beta\in\bar{\Delta}$ is non-regular, so has a conjugate not involving a root element for $\beta$ (by [22, Tab. 4]) and hence we are reduced to the previous case since ${\mathbf{C}}_{\mathbf{G}}^{\circ}(\sigma)$ has rank at least 3. Note that this also holds trivially if this Levi is of type $A_{2}^{2}$. ∎
@@ 6.3. Subnormalisers of semisimple elements
What do subnormalisers of semisimple elements look like in connected reductive groups? They are related to the $p$-closed subsystems of $\Phi$ (see [24, Def. 13.2]); as before it is easy and convenient to reduce to the case of simple groups.
@@
Let $s\in{\mathbf{G}}$ be semisimple and ${\mathbf{T}}$ a maximal torus of ${\mathbf{G}}$ containing $s$ (which exists by [24, Cor. 6.11]). Then ${\mathbf{N}}_{\mathbf{G}}({\mathbf{T}})$ is contained in ${\operatorname{Sub}}_{\mathbf{G}}(s)$, so ${\mathbf{T}}$ normalises ${\operatorname{Sub}}_{\mathbf{G}}(s)$. Let $\Psi:=\{\alpha\in\Phi\mid{\mathbf{U}}_{\alpha}\leq{\operatorname{Sub}}_{\mathbf{G}}(s)\}$. Then, $\Psi$ is a $p$-closed subsystem of $\Phi$, by definition. Since ${\operatorname{Sub}}_{\mathbf{G}}(s)$ contains ${\mathbf{N}}_{\mathbf{G}}({\mathbf{T}})$ and hence representatives for all Weyl group elements, $\Psi$ consists of full $W$-orbits of roots, that is, of all short, all long, or all roots, or $\Psi=\emptyset$ (see [24, Cor. A.18]). Now ${\operatorname{Sub}}_{\mathbf{G}}(s)^{\circ}$ is a subsystem subgroup since it is normalised by the maximal torus ${\mathbf{T}}$. According to [24, Cor. 13.7] this shows that ${\operatorname{Sub}}_{\mathbf{G}}(s)={\mathbf{N}}_{\mathbf{G}}({\mathbf{G}}(\Psi))$. ∎
@@
By Theorem 6.8 we have ${\operatorname{Sub}}_{\mathbf{G}}(s)={\mathbf{N}}_{\mathbf{G}}({\mathbf{G}}(\Psi))$ for some $p$-closed subsystem $\Psi$ of $\Phi$ consisting of full orbits of roots under the Weyl group. If $\Psi=\emptyset$ then ${\mathbf{G}}(\Psi)={\mathbf{T}}$ and thus ${\operatorname{Sub}}_{\mathbf{G}}(s)={\mathbf{N}}_{\mathbf{G}}({\mathbf{T}})$. For the rest of the proof assume $\Psi\neq\emptyset$. We need to understand the possible $\Psi$, where, of course, we may assume $\Phi$ has two root lengths. The structure of such subsystems is given in [24, Tab. B.2]. By [24, Thm 13.14 and Prop. 13.15] these are $p$-closed precisely under the conditions listed in the statement. Note that all of these subgroups do indeed contain the full normaliser of a maximal torus of ${\mathbf{G}}$.

### v2

@@ 6.1. Subnormalisers of unipotent elements
By [27, Prop. 2.9] we may embed $\mathbf{U}$} into the (unipotent) group of upper uni-triangular matrices of ${\operatorname{GL}}_{n}$ for a suitable $n$, so without loss we may assume $\mathbf{U}$} is connected. By [27, Cor. 2.10] the unipotent group $\mathbf{U}$} is nilpotent, so by [19, Prop. 17.4], for any proper closed subgroup ${\mathbf{H}}<{\mathbf{U}}$ we have $\dim{\mathbf{H}}<\dim{\mathbf{N}}_{\mathbf{U}}({\mathbf{H}})$. Thus by induction on the dimension, any closed subgroup is subnormal in $\mathbf{U}$}. As $\langle x\rangle$ is normal in its abelian (see [19, Lemma 15.1.C]) closure $\overline{\langle x\rangle}$ this achieves the proof. ∎
@@
Now the Zariski closure $\overline{{\widehat{R_{u}}}(H)}$ is unipotent [27, Prop. 2.9], giving $\langle x\rangle{\,\triangleleft\triangleleft\,}\overline{{\widehat{R_{u}}}(H)}$ by Lemma 6.1 and then of course $\langle x\rangle{\,\triangleleft\triangleleft\,}{\widehat{R_{u}}}(H)$. Hence, $\langle x\rangle{\,\triangleleft\triangleleft\,}{\mathbf{N}}_{\mathbf{G}}({\widehat{R_{u}}}(H))\geq H$ and the last assertion follows. ∎
@@
Let ${\mathbf{U}}\leq{\mathbf{G}}$ be a closed unipotent subgroup with ${\mathbf{U}}={\widehat{R_{u}}}({\mathbf{N}}_{\mathbf{G}}({\mathbf{U}}))$. By [27, Cor. 17.15], $\mathbf{U}$} lies in some Borel subgroup of $\mathbf{G}$} and then by [27, Thm 17.10] there is a parabolic subgroup ${\mathbf{P}}\leq{\mathbf{G}}$ with ${\mathbf{U}}\leq R_{u}({\mathbf{P}})$ and ${\mathbf{N}}_{\mathbf{G}}({\mathbf{U}})\leq{\mathbf{P}}$. In particular, ${\mathbf{N}}_{\mathbf{G}}({\mathbf{U}})$ normalises $R_{u}({\mathbf{P}})$. Thus, $V:={\mathbf{N}}_{R_{u}({\mathbf{P}})}({\mathbf{U}})$ is contained in and normalised by ${\mathbf{N}}_{\mathbf{G}}({\mathbf{U}})$, so contained in ${\widehat{R_{u}}}({\mathbf{N}}_{\mathbf{G}}({\mathbf{U}}))={\mathbf{U}}$, whence $V={\mathbf{U}}$. By [19, Prop. 17.4] this forces ${\mathbf{U}}=R_{u}({\mathbf{P}})$.
Conversely, if ${\mathbf{P}}\leq{\mathbf{G}}$ is parabolic, hence connected, then ${\mathbf{P}}/R_{u}({\mathbf{P}})$ is connected reductive, and so has no non-trivial normal unipotent subgroups (since any such would have to be finite, hence central, but all central elements of connected reductive groups are semisimple, e.g. by [27, Cor. 8.13(b)]), whence ${\widehat{R_{u}}}({\mathbf{P}})=R_{u}({\mathbf{P}})$. ∎
@@
By Lemma 6.1, ${\operatorname{Sub}}_{\mathbf{G}}(x)$ contains all Borel subgroups of $\mathbf{G}$} containing $x$. For the converse, assume $\langle x\rangle{\,\triangleleft\triangleleft\,}H$ for some subgroup $H\leq{\mathbf{G}}$, where without loss of generality $H$ is maximal with respect to inclusion. By Lemma 6.2 then $H={\mathbf{N}}_{\mathbf{G}}({\widehat{R_{u}}}(H))$ and $x\in Q:={\widehat{R_{u}}}(H)$. Now $Q$ lies in the unipotent radical $\mathbf{U}$} of some Borel subgroup of $\mathbf{G}$} ([27, Cor. 17.15]). If $Q<{\mathbf{U}}$ then $Q_{1}:={\mathbf{N}}_{\mathbf{U}}(Q)>Q$, and $\langle x\rangle{\,\triangleleft\triangleleft\,}\langle H,Q_{1}\rangle>H$, a contradiction, so in fact $Q={\mathbf{U}}$ and $H={\mathbf{N}}_{\mathbf{G}}(Q)={\mathbf{N}}_{\mathbf{G}}({\mathbf{U}})$ is closed (see [27, Ex. 10.18]). Then ${\mathbf{N}}_{H}({\mathbf{U}})^{\circ}$ is a Borel subgroup of $H$, and thus $H^{\circ}$ is generated by the $H$-conjugates of ${\mathbf{N}}_{H}({\mathbf{U}})^{\circ}$ (see [27, Thm 6.10]), all of which contain $Q$ and hence $x$. Thus, $H$ is generated by the normalisers (in $H$) of the maximal unipotent subgroups (of $H$ and hence of $G$) it contains. The claim follows. ∎
@@ 6.2. An extension to disconnected groups
In this subsection we consider a slightly different situation. Let now $\mathbf{G}$} be the extension of a connected reductive algebraic group $\mathbf{G}$$^{\circ}$ in characteristic $p$ by a graph automorphism $\sigma$ of order $p$. We are interested in unipotent elements in non-trivial cosets of $\mathbf{G}$$^{\circ}$ in $\mathbf{G}$}. Every coset of $\mathbf{G}$$^{\circ}$ in $\mathbf{G}$} contains a unique class of regular unipotent elements satisfying analogous properties to the case of connected groups (see [36, Prop. II.10.2]).
Let ${\mathbf{B}}\leq{\mathbf{G}}$ be the normaliser in $\mathbf{G}$} of a $\sigma$-stable Borel subgroup $\mathbf{B}$$^{\circ}$ of $\mathbf{G}$$^{\circ}$, with $\sigma$-stable maximal torus $\mathbf{T}$$^{\circ}$ and unipotent radical ${\mathbf{U}}={\mathbf{U}}^{\circ}$. Let $W={\mathbf{N}}_{{\mathbf{G}}^{\circ}}({\mathbf{T}}^{\circ})/{\mathbf{T}}^{\circ}$ be the Weyl group of $\mathbf{G}$$^{\circ}$; it is normalised by $\sigma$, and $\sigma$ acts on its set of roots $\Phi$ with respect to $\mathbf{T}$$^{\circ}$, permuting its set $\Delta$ of simple roots. We let $\bar{\Delta}$ denote the set of $\sigma$-orbits in $\Delta$, and we write $\bar{S}$ for the set of simple reflections of $W^{\sigma}$ constructed as in [27, Lemma 23.3].
@@
Since $\mathbf{U}$} is a maximal unipotent subgroup in $\mathbf{G}$$^{\circ}$ and normalised by $\sigma$ we may assume $x\in{\mathbf{U}}\sigma$. So $x=x_{0}\sigma$ where $x_{0}$ is a product of root elements for $\mathbf{G}$$^{\circ}$. If root elements for representatives from all $\sigma$-orbits of simple roots occur in $x_{0}$, then $x$ is regular and hence contained in a unique Borel subgroup of $\mathbf{G}$} by [36, Prop. II.10.2], so ${\operatorname{Sub}}_{\mathbf{G}}(x)={\mathbf{C}}_{\mathbf{B}}(\sigma)$. Representatives for the non-regular outer unipotent classes in types $A_{2},A_{3}$ and $A_{4}$ are given in [24, Tab. 4]. It transpires that in each case at least one of the given representatives in the finite group for a class in the algebraic group has the property that its image under any reflection in $\bar{S}$ is still in ${\mathbf{U}}\sigma$, and thus $x$ lies in a unipotent normal subgroup of any ${\mathbf{P}}^{\sigma}$, where $\mathbf{P}$} runs over the normalisers in $\mathbf{G}$} of the minimal $\sigma$-stable standard parabolic subgroups of $\mathbf{G}$$^{\circ}$. Since these generate $\mathbf{G}$}, we have ${\operatorname{Sub}}_{\mathbf{G}}(x)={\mathbf{G}}$. ∎
@@
First, by [36, Prop. II.10.2] the regular unipotent elements in ${\mathbf{G}}^{\circ}\sigma$ are picky, arguing as in the proof of Proposition 3.1. Now assume $x\in{\mathbf{G}}^{\circ}\sigma$ is unipotent non-regular. Again we may assume $x\in{\mathbf{U}}\sigma$. If ${\mathbf{C}}_{{\mathbf{G}}^{\circ}}(\sigma)$ has rank 1 or 2, we are done by Proposition 6.6, or $\mathbf{G}$$^{\circ}$ is of type $D_{4}$, $p=3$ and $\sigma$ induces triality. In the latter case, inspection of the representatives given in [24, Tab. 8] shows that $x$ can be chosen to lie in the maximal normal unipotent subgroups of ${\mathbf{C}}_{\mathbf{P}}(\sigma)$, where $\mathbf{P}$} runs over the normalisers in $\mathbf{G}$} of the minimal $\sigma$-stable standard parabolic subgroups of $\mathbf{G}$$^{\circ}$ and thus ${\operatorname{Sub}}_{\mathbf{G}}(x)={\mathbf{G}}$.
So we may now assume that $p=2$, ${\mathbf{C}}_{\mathbf{G}}^{\circ}(\sigma)$ has rank at least 3, and that the claim has been shown for all groups of smaller rank. As $x$ is not regular, by [36, Prop. II.10.2] it is a product of root elements not involving representatives from at least one orbit of simple roots $\alpha\in\bar{\Delta}$. Consider the Dynkin diagram of ${\mathbf{C}}_{{\mathbf{G}}}^{\circ}(\sigma)$, with nodes labelled by $\bar{\Delta}$. If $\alpha$ is not an end node, then we can conclude verbatim by the arguments in the proof of Proposition 3.6. If $\alpha$ is an end node, then the image of $x$ in the standard Levi subgroup corresponding to $\alpha$ and the adjacent node $\beta\in\bar{\Delta}$ is non-regular, so has a conjugate not involving a root element for $\beta$ (by [24, Tab. 4]) and hence we are reduced to the previous case since ${\mathbf{C}}_{\mathbf{G}}^{\circ}(\sigma)$ has rank at least 3. Note that this also holds trivially if this Levi is of type $A_{2}^{2}$. ∎
@@ 6.3. Subnormalisers of semisimple elements
What do subnormalisers of semisimple elements look like in connected reductive groups? They are related to the $p$-closed subsystems of $\Phi$ (see [27, Def. 13.2]); as before it is easy and convenient to reduce to the case of simple groups.
@@
Let $s\in{\mathbf{G}}$ be semisimple and $\mathbf{T}$} a maximal torus of $\mathbf{G}$} containing $s$ (which exists by [27, Cor. 6.11]). Then ${\mathbf{N}}_{\mathbf{G}}({\mathbf{T}})$ is contained in ${\operatorname{Sub}}_{\mathbf{G}}(s)$, so $\mathbf{T}$} normalises ${\operatorname{Sub}}_{\mathbf{G}}(s)$. Let $\Psi:=\{\alpha\in\Phi\mid{\mathbf{U}}_{\alpha}\leq{\operatorname{Sub}}_{\mathbf{G}}(s)\}$. Then, $\Psi$ is a $p$-closed subsystem of $\Phi$, by definition. Since ${\operatorname{Sub}}_{\mathbf{G}}(s)$ contains ${\mathbf{N}}_{\mathbf{G}}({\mathbf{T}})$ and hence representatives for all Weyl group elements, $\Psi$ consists of full $W$-orbits of roots, that is, of all short, all long, or all roots, or $\Psi=\emptyset$ (see [27, Cor. A.18]). Now ${\operatorname{Sub}}_{\mathbf{G}}(s)^{\circ}$ is a subsystem subgroup since it is normalised by the maximal torus $\mathbf{T}$}. According to [27, Cor. 13.7] this shows that ${\operatorname{Sub}}_{\mathbf{G}}(s)={\mathbf{N}}_{\mathbf{G}}({\mathbf{G}}(\Psi))$. ∎
@@
By Theorem 6.8 we have ${\operatorname{Sub}}_{\mathbf{G}}(s)={\mathbf{N}}_{\mathbf{G}}({\mathbf{G}}(\Psi))$ for some $p$-closed subsystem $\Psi$ of $\Phi$ consisting of full orbits of roots under the Weyl group. If $\Psi=\emptyset$ then ${\mathbf{G}}(\Psi)={\mathbf{T}}$ and thus ${\operatorname{Sub}}_{\mathbf{G}}(s)={\mathbf{N}}_{\mathbf{G}}({\mathbf{T}})$. For the rest of the proof assume $\Psi\neq\emptyset$. We need to understand the possible $\Psi$, where, of course, we may assume $\Phi$ has two root lengths. The structure of such subsystems is given in [27, Tab. B.2]. By [27, Thm 13.14 and Prop. 13.15] these are $p$-closed precisely under the conditions listed in the statement. Note that all of these subgroups do indeed contain the full normaliser of a maximal torus of $\mathbf{G}$}.